In a particular town there are 100 cars and 50 motorcycles. Each car owner has a demand curve for petrol given by: Qdc = 20 – 5p for p ≤ 4 and Qdc = 0 for p > 4. Each motorcycle owner has the following demand function for petrol: Qdm = 15 – 3p for p ≤ 5 and Qdm = 0 for p > 5. Prices are expressed as SR per gallon, and quantities as gallons per week. (5marks)
1
Expert's answer
2017-11-14T14:59:07-0500
In a particular town there are 100 cars and 50 motorcycles. Each car owner has a demand curve for petrol given by: Qdc = 20 – 5p for p ≤ 4 and Qdc = 0 for p > 4. Each motorcycle owner has the following demand function for petrol: Qdm = 15 – 3p for p ≤ 5 and Qdm = 0 for p > 5.
a) If the price is SR 3, then:
(i) each car owner will buy Q = 20 - 5*3 = 5 gallons of petrol per week.
(ii) each motorcycle owner will buy Q = 15 - 3*3 = 6 gallons of petrol per week.